the hypotenuse in a right-angled isosceles triangle measures 4.5cm. determine the length of the other two sides

A= 3.18cm

Answer by stanbon(75887)

the hypotenuse in a right-angled isosceles triangle measures 4.5cm. determine the length of the other two sides
A= 3.18cm
-----
Let the length of the 2 equal sides each be "x".
Equation;
x^2 + x^2 = 4.5^2
2x^2 = 20.25
x^2 = 10.125
x = 3.18cm
===============
Cheers,
Stan H.